3.5 Multiply 62 by 35, firstly by breaking down 35 in its terms (30+5) and secondly by breaking down 35 in its factors (5×7). Show all your steps


1
Expert's answer
2022-06-02T18:23:35-0400

a)


"62(35)=62(30+5)=62(30)+62(5)"

"=1860+310=2170"

b)


"62(35)=62(5\times 7)=62(5)\times 7"

"=310\times 7=2170"